# On the Performance of the Neyman Allocation with Small Pilots

### v3

@@ 4.4 Simulation Evidence
The efficiency losses recorded in Proposition 4.7 are extremely pessimistic and are likely not relevant in practice. To better appreciate the losses that researchers might experience, we next present the mean-squared error for $\hat{\theta}_{\tilde{p}}$ under various data-generating processes, computed via Monte Carlo simulations. Our choice of parameter values are inspired by [6, ].
Let $Y(1)=0.075+\sigma(1)\varepsilon(1)$ and $Y(0)=\sigma(0)\varepsilon(0)$. We are interested in how the MSE of $\hat{\theta}_{\tilde{p}}$ as $\sigma(1)/\sigma(0)$ varies over $[0,1]$. We do not need to consider values greater than $1$ given the symmetry. We set $\sigma^{2}(1)+\sigma^{2}(0)=2$ so that the variance of $\hat{\theta}_{\frac{1}{2}}$ is fixed. As in Section 3, consider the following models:
@@
Next, consider Model 1. With a normal distribution, the MSE of the FNA is 2.9% larger than the balanced allocation when $\sigma(1)/\sigma(0)=1$. Weighed against a potential reduction of 50%, the FNA is definitely preferred in this setting. However, the picture becomes less clear once we move to model 2. Here, the MSE of the FNA is around 17% larger than that of the balanced allocation. This is true up to $\sigma(1)/\sigma(0)=0.75$, which is more heteroskedasticity in many of the outcomes considered in our empirical examples. Similar results obtain with Models 3,4 and 5, in which FNA has MSE that are 26%, 20% and 17% more that of the balanced allocation when $\sigma(1)/\sigma(0)=1$. The above results suggests that the FNA may not be advisable when $m=20$. On the other hand, the worst-case efficiency loss shrinks quickly as $m$ increases, so that the FNA performs favourably. Results for $m=50$ can be found in Appendix C.1. At $m=50$, the loss in efficiency over models 1 to 5 are, respectively, 1.5%, 6.4%, 13%, 11% and 8.6%.
Figure 3: MSE under the Infeasible Neyman Allocation (INA), Balanced Allocation (BA) and Feasible Neyman Allocation (FNA) when $\varepsilon(1),\varepsilon(0)$ has the standard Normal distribution and the standardized Chi-Squared distribution.
@@ 5 Empirical Evidence of Approximate Homoskedasticity
To assess the amount of heteroskedasticity that empirical researchers face, we revisit the first 10 completed experiments in the AER RCT Registry. In each experiment, we ask the following question: Suppose the authors had access to a small pilot prior to the main study, would they have done better using the FNA instead of balanced randomization? In each experiment presented, we use the full experimental sample to estimate the standard deviations of each treatment arm and compute the corresponding ratios to see if these are close to one. In practice, researchers cannot do this given a small pilot since they do not have access to consistent variance estimators. Our findings suggest that these authors would likely not have not better with the FNA. We present two examples in this section: [2, ] is an experiment in which the outcomes are relatively homoskedastic. [1, ] contains outcome variables which are heteroskedastic. In this example, we also provide estimates of the interval $C_{m}$ and show that it will be wide even when $m$ is large. The remaining eight experiments are qualitatively similar to [2, ] and can be found in Section B of the online appendix.
5.1 [2, ]
@@
There is a significant body of research examining the impact of school-level factors such as class size or teacher quality on educational performance of students. These are typically seen as the primary instruments for educational policy intervention. A large body of work also examines the impact of parental inputs on educational outcomes. [2, ] study whether or not parental inputs can be effectively manipulated through simple participation programs at schools. They do so via a large-scale randomized control trial in middle schools in the Créteil educational district of Paris. The experiment targeted families of 6th graders and the program consisted of a sequence of three meetings with parents every 2–3 weeks. The sessions focused on how parents can help their children by participating at school and at home in their education and additionally, included advice on how to adapt to results in end-of-term report cards. Participation in the program was randomized at the class level – half of the classes at a given school were assigned to the participation program. Classes are groups of 20–30 students. The overall sample comprised of 183 classes and a total of 4,308 students. The study tracked three types of outcomes: (1) parental involvement attitudes and behaviour; (2) children’s behaviour, namely truancy, disciplinary record and work effort; and (3) children’s academic results. Since randomization was done at the class-level, we examine heteroskedasticity with respect to treatment status at both the individual level and at the class level.
@@
Table 1 reports student-level standard deviations in treatment and control groups, as well as their ratios, for the main outcomes of interest. These are the outcomes considered in Tables 2, 3 and 5 in [2, ]. The ratios are all close to one, so that by and large the treatment and control groups are relatively homoskedastic at the student level. This indicates that if the experimenters were to run a randomized control trial in the same population with treatment assigned at the individual-level, the FNA would likely yield no improvement relative to the balanced allocation.
@@
Table 2 reports standard deviations and their ratios in class-level means of outcomes. This corresponds to a scenario in which classes are the units of interest, with class-level means as the relevant outcomes. We first calculate class-level means and then compute their standard deviations across classes for the treatment and control groups respectively. The standard deviation ratios for class-level means are by and large also close to one. Hence, if the hypothetical experiment was to be conducted at the class-level, the treatment and control groups would still be relatively homoskedastic. In this case, the FNA would again not improve upon the balanced allocation.
Table 1: Student-Level Heteroskedasticity in [2, ].
@@
Table 2: Class-Level Heteroskedasticity in [2, ].
@@
5.2 [1, ]
A large body of economic models posit that individuals have time inconsistent preferences, exhibiting more impatience for near-term trade-offs than for future trade-offs. The implication of these models is that those who engage in commitment devices ex ante may improve their welfare. To test this hypothesis, [1, ] conducted an RCT in the Philippines, in which individuals were offered randomly offered the chance to open a SEED (Save, Earn, Enjoy Deposits) account. Money deposited into the account cannot be withdrawn until the owner reached a goal, such as reaching a savings amount or until a pre-specified month in which they expected large expenditures.
Partnering with a rural bank in Mindanao, they authors surveyed 1,777 of their existing or former clients, of which 842 were placed into the treatment group, while 469 were placed in the control group. As treatment involved receiving a briefing on the importance of savings, the remaining 466 individuals were placed in the marketing group, receiving the briefing but not access to SEED. We focus on the Table VI of [1, ], containing results on saving behavior. The parameter of interest is the Intent-to-Treat effect, with approximately 25% of the treated taking up treatment. Here, the authors find that relative to the control group, the treatment group had a higher change in savings 6 months (6m) and 12 months (12m) after treatment. Comparing treatment to marketing group led to weaker but still positive results.
@@
We first note that $\Delta$ Tot. Bal. $>0\%$ (12m) and $\Delta$ Tot. Bal. $>20\%$ (12m) are binary outcomes which are relatively homoskedastic. $\Delta$ Tot. Bal. (6m) and $\Delta$ Tot. Bal. (12m), measured in Philippine pesos, exhibit more heteroskedasticity. In particular, comparing the treatment group to the marketing group at the 12 month period, we observe a standard deviation ratio of $3.13$. At first glance, this suggests that the FNA might outperform balanced randomization, at least with respect to this specific outcome. This turns out to be false once we investigate the heteroskedasticity in $\Delta$ Tot. Bal. (6m) and $\Delta$ Tot. Bal. (12m). Quantiles of these variables are displayed in Table 4. Clearly, they have extremely fat right tails, which we confirm by computing the kurtosis, contained in Table 5. Fat tails worsen the performance of a variety of statistical techniques, including the FNA, as our analysis in Section 4 shows.
Table 3: Heteroskedasticity in [1, ]
@@
Table 4: Quantiles of Outcome Variables in [1, ].
@@
Table 5: Kurtosis of the Outcome Variables in [1, ]
@@
Researchers cannot estimate $C_{m}$ if they only have access to a small pilot. However, using data from the full experiment, we are able to estimate $C_{m}$ for a range of $m$. To do so, we draw with replacement samples of size $m$ and compute $Z_{m}$ for each draw. We then take the average of $Z_{m}+1/Z_{m}$ over 10,000 draws to evaluate $B_{m}$, essentially evaluating the expectation with respect to the empirical CDF. Proposition 4.4 then allows us to compute $C_{m}$. We evaluate $C_{m}$ for $m\leq 2000$, using grids of size $20$. Results are displayed in Figure 5. Given the high kurtosis, there is a relatively large range of ratios of standard deviations for which the balanced allocation is preferred to the FNA.
We can further compute the $m$ necessary before the FNA outperforms the balanced allocation. Given the upper and lower bounds in Figure 5, we smooth over the grid points using cubic interpolation. We then find the minimum $m$ at which the either bound attains the values in Table 3. These are the “Exact” intervals presented in Table 6. Here, we see that the necessary pilot sizes are between 25–50% of the full experiment. For the comparison of Treatment and Marketing group at 6 months, the necessary pilot size exceeds $2,000$, falling outside the set of grid points we explored. To complete the analysis, we use Proposition 4.5 to obtain approximations of the necessary $m$. Specifically, we estimate $V$ using their sample analogs computed on the full experiment:
$\small\hat{V}=\frac{\frac{1}{n_{1}}\sum_{A_{i}=1}\left(Y_{i}-\overline{Y}(1)\right)^{4}}{\left(\frac{1}{n_{1}-1}\sum_{A_{i}=1}\left(Y_{i}-\overline{Y}(1)\right)^{4}\right)^{2}}+\frac{\frac{1}{n_{0}}\sum_{A_{i}=0}\left(Y_{i}-\overline{Y}(0)\right)^{4}}{\left(\frac{1}{n_{0}-1}\sum_{A_{i}=0}\left(Y_{i}-\overline{Y}(0)\right)^{4}\right)^{2}}-2\quad,\quad\overline{Y}(a)=\frac{1}{n_{a}}\sum_{A_{i}=a}Y_{i}$
where $n_{a}$ is the number of units with treatment status $a$ in the full experiment. We can then estimate $m$ by
$\hat{m}=\frac{\hat{V}}{\left(1-\frac{\hat{\sigma}(1)}{\hat{\sigma}(0)}\right)^{2}}$
Comparing the asymptotic intervals to the exact ones, we see that the former is far too optimistic for the fat-tailed DGP in [1, ]. This is likely because the sub-Gaussian assumption is inappropriate for this data. Nonetheless, applying the asymptotic bounds to the case of Treatment vs Marketing group at 6 months, we find that a pilot size of 7,000 would be necessary for the FNA to outperform balanced randomization. This is nearly 4 times the size of the actual experiment.
@@
In sum, our analysis suggests that the FNA would have performed poorly in the context of [1, ]. Even though the outcomes of interest exhibit stronger heteroskedasticity, the fat tails of the outcome distributions also impedes the estimation of $\widetilde{p}$, so that ultimately, very large pilot sizes are needed for the FNA to outperform the balanced allocation in this example.
@@ 6 Potential Solutions
In this section, we explore, using simulations methods that can improve the performance of the FNA with small pilots. For definition of the models, see Section 4.4. We consider in turn (1) testing for homoskedasticity, (2) additive regularization and (3) exponential regularization. We conclude with a comparison of these results and speculate that exponential regularization is preferable to the other two methods. Additional simulation results can be found in Appendix C.2.
@@ 6.1 Testing for Homoskedasticity
Since balanced randomization is preferred when outcome are relatively homoskedastic, consider the following two-step procedure: First test for homoskedasticity in the pilot. If the test rejects, use the FNA in the main wave. Otherwise, use the balanced allocation instead.
@@
$\tilde{W}=\sqrt{\frac{m}{2}}\frac{(\tilde{\sigma}^{2}(1)-\tilde{\sigma}^{2}(0))}{\sqrt{\tilde{V}_{W}(1)+\tilde{V}_{W}(0)}}\leavevmode\nobreak\ ,$
@@
$\tilde{V}_{W}(a)=\begin{pmatrix}1&-2\overline{Y}(a)\end{pmatrix}\begin{pmatrix}\frac{1}{m/2}\sum_{A_{i}=a}(Y^{2}_{i}-\overline{Y^{2}}(a))^{2}&\frac{1}{m/2}\sum_{A_{i}=a}(Y_{i}-\overline{Y}(a))(Y^{2}_{i}-\overline{Y^{2}}(a))\\ *&\frac{1}{m/2}\sum_{A_{i}=a}(Y_{i}-\overline{Y}(a))^{2}\end{pmatrix}\begin{pmatrix}1\\ -2\overline{Y}(a)\end{pmatrix}\leavevmode\nobreak\ ,$
$\overline{Y}(a)=\frac{1}{m/2}\sum_{A_{i}=a}Y_{i}\quad,\quad\overline{Y^{2}}(a)=\frac{1}{m/2}\sum_{A_{i}=a}Y^{2}_{i}\leavevmode\nobreak\ .$
@@
$\phi=\mathbf{1}\left\{\tilde{W}>\Phi^{-1}\left(1-\frac{\alpha}{2}\right)\right\}+\mathbf{1}\left\{\tilde{W}<\Phi^{-1}\left(\frac{\alpha}{2}\right)\right\}\leavevmode\nobreak\ ,$
where $\Phi(\cdot)$ is the CDF of the standard normal distribution. Under small pilot asymptotics, this test will not be consistent. However, the two-step procedure will be asymptotically equivalent to the FNA when pilots are large, since for any fixed level $\alpha$, the test will reject with probability approaching $1$ as $m\to\infty$. However, under such a “pre-testing procedure”, the main-wave difference-in-means estimator will not be uniformly consistent.
Nonetheless, the procedure appears to perform well in practice. Figure 7 present the MSEs that obtain using first stage tests with $\alpha\in\left\{1\%,5\%,10\%,20\%,50\%\right\}$. When $\sigma(1)/\sigma(0)\geq 0.5$, decreasing $\alpha$ decreases MSE across all distributions. Conversely, when $\sigma(1)/\sigma(0)\leq 0.5$, decreasing $\alpha$ increases MSE across all distributions. The appropriate level of $\alpha$ then depends on researcher’s preference for the performance of the method over the range of $\sigma(1)/\sigma(0)$
Figure 6: MSE under the Infeasible Neyman Allocation (INA), Balanced Allocation (BA) and Feasible Neyman Allocation (FNA) when $\varepsilon(1),\varepsilon(0)$ has the standard Normal distribution and the standardized Chi-Squared distribution.
Figure 6: MSE under the Infeasible Neyman Allocation (INA), Balanced Allocation (BA) and Feasible Neyman Allocation (FNA) when $\varepsilon(1),\varepsilon(0)$ has the standard Normal distribution and the standardized Chi-Squared distribution.
Figure 7: MSE under the Infeasible Neyman Allocation (INA), Balanced Allocation (BA) and Feasible Neyman Allocation (FNA) when $\varepsilon(1),\varepsilon(0)$ has the standard Normal distribution and the standardized Chi-Squared distribution.

### v4

@@ 4.4 Simulation Evidence
The efficiency losses recorded in Proposition 4.7 are pessimistic and may not be relevant in practice. To better appreciate the losses that researchers might experience, we next present mean-squared errors for $\hat{\theta}_{\tilde{p}}$ under various data-generating processes, computed via Monte Carlo simulations. Our choice of parameter values is inspired by Blackwell et al. (2022).
Let $Y(1)=0.075+\sigma(1)\varepsilon(1)$ and $Y(0)=\sigma(0)\varepsilon(0)$. We are interested in the MSE of $\hat{\theta}_{\tilde{p}}$ as $\sigma(1)/\sigma(0)$ varies over $[0,1]$. We do not need to consider values greater than $1$ given the reciprocal symmetry. We set $\sigma^{2}(1)+\sigma^{2}(0)=2$ so that the variance of $\hat{\theta}_{\frac{1}{2}}$ is fixed. As in Section 3, consider the following models:
@@
Next, consider Model 1. With a normal distribution, the MSE of the FNA is 2.9% larger than the balanced allocation when $\sigma(1)/\sigma(0)=1$. Weighed against a potential reduction of 50%, the FNA is definitely preferred in this setting. However, the picture becomes less clear once we move to Model 2. Here, the MSE of the FNA is around 17% larger than that of the balanced allocation. This is true up to $\sigma(1)/\sigma(0)=0.75$, which is more heteroskedasticity in many of the outcomes considered in our empirical examples. Similar results obtain with Models 3, 4 and 5, in which the FNA has MSEs that are 26%, 20% and 17% more that of the balanced allocation when $\sigma(1)/\sigma(0)=1$. The above results suggest that the FNA may not be advisable when $m=20$. On the other hand, the worst-case efficiency loss shrinks quickly as $m$ increases, so that the FNA performs favorably. Results for $m=50$ can be found in Appendix C.1. At $m=50$, the loss in efficiency over Models 1 to 5 are, respectively, 1.5%, 6.4%, 13%, 11% and 8.6%.
Figure 3: MSE under the Infeasible Neyman Allocation (INA), Balanced Allocation (BA) and Feasible Neyman Allocation (FNA) when $\varepsilon(1),\varepsilon(0)$ has the standard Normal distribution and the standardized Chi-Squared distribution. $m=20$.
Figure 4: MSE under the Infeasible Neyman Allocation (INA), Balanced Allocation (BA) and Feasible Neyman Allocation (FNA) when $\varepsilon(1),\varepsilon(0)$ has the standardized Pareto($1,s$) distribution, with $s\in\{3,4,5\}$. $m=20$.
@@ 5 Empirical Evidence of Approximate Homoskedasticity
To assess the amount of heteroskedasticity that empirical researchers face, we revisit the first 10 completed experiments in the AER RCT Registry. In each experiment, we ask the following question: suppose the authors had access to a small pilot prior to the main study, would they have done better using the FNA instead of balanced randomization? In each experiment presented, we use the full experimental sample to estimate the standard deviations of each treatment arm and compute the corresponding ratios to see if these are close to one. In practice, researchers cannot do this given a small pilot since they do not have access to consistent variance estimators. Our findings suggest that these authors would likely not have done better with the FNA. We present two examples in this section: Avvisati et al. (2014) conduct an experiment in which the outcomes are relatively homoskedastic. Ashraf et al. (2006) study outcome variables which are heteroskedastic. In this latter example, we also provide estimates of the interval $C_{m}$ and show that it will be wide even when $m$ is large. The remaining eight experiments are qualitatively similar to Avvisati et al. (2014) and can be found in Section B of the online appendix.
5.1 Avvisati et al. (2014)
@@
There is a significant body of research examining the impact of school-level factors such as class size or teacher quality on educational performance of students. These are typically seen as the primary instruments for educational policy intervention. A large body of work also examines the impact of parental inputs on educational outcomes. Avvisati et al. (2014) study whether or not parental inputs can be effectively manipulated through simple participation programs at schools. They do so via a large-scale randomized control trial in middle schools in the Créteil educational district of Paris. The experiment targeted families of 6th graders and the program consisted of a sequence of three meetings with parents every 2–3 weeks. The sessions focused on how parents can help their children by participating at school and at home in their education and additionally included advice on how to adapt to results in end-of-term report cards. Participation in the program was randomized at the class level – half of the classes at a given school were assigned to the participation program. Classes were groups of 20–30 students. The overall sample comprised 183 classes and a total of 4,308 students. The study tracked three types of outcomes: (1) parental involvement attitudes and behavior; (2) children’s behavior, namely truancy, disciplinary record and work effort; and (3) children’s academic results. Since randomization was done at the class-level, we examine heteroskedasticity with respect to treatment status at both the individual level and at the class level.
@@
Table 1 reports student-level standard deviations in treatment and control groups, as well as their ratios, for the main outcomes of interest. These are the outcomes considered in Tables 2, 3 and 5 in Avvisati et al. (2014). The ratios are all close to one, so that by and large the treatment and control groups are relatively homoskedastic at the student level. This indicates that if the experimenters were to run a randomized control trial in the same population with treatment assigned at the individual-level, the FNA would likely yield no improvement relative to the balanced allocation.
@@
Table 2 reports standard deviations and their ratios in class-level means of outcomes. This corresponds to a scenario in which classes are the units of interest, with class-level means as the relevant outcomes. We first calculate class-level means and then compute their standard deviations across classes for the treatment and control groups respectively. The standard deviation ratios for class-level means are by and large also close to one. Hence, if the hypothetical experiment were to be conducted at the class-level, the treatment and control groups would still be relatively homoskedastic. In this case, the FNA would again not improve upon the balanced allocation.
Table 1: Student-Level Heteroskedasticity in Avvisati et al. (2014).
@@
Table 2: Class-Level Heteroskedasticity in Avvisati et al. (2014).
@@
5.2 Ashraf et al. (2006)
A large body of economic models posit that individuals have time inconsistent preferences, exhibiting more impatience for near-term trade-offs than for future trade-offs. The implication of these models is that those who engage in commitment devices ex ante may improve their welfare. To test this hypothesis, Ashraf et al. (2006) conducted an RCT in the Philippines, in which individuals were offered randomly offered the chance to open a SEED (Save, Earn, Enjoy Deposits) account. Money deposited into the account cannot be withdrawn until the owner reached a goal, such as reaching a savings amount or until a pre-specified month in which they expected large expenditures.
Partnering with a rural bank in Mindanao, the authors surveyed 1,777 of their existing or former clients, of which 842 were placed into the treatment group, while 469 were placed in the control group. As treatment involved receiving a briefing on the importance of savings, the remaining 466 individuals were placed in the marketing group, receiving the briefing but not access to SEED. We focus on the Table VI of Ashraf et al. (2006), containing results on saving behavior. The parameter of interest is the Intent-to-Treat effect, with approximately 25% of the treated taking up treatment. Here, the authors find that relative to the control group, the treatment group had a higher change in savings 6 months (6m) and 12 months (12m) after treatment. Comparing the treatment to the marketing group led to weaker but still positive results.
@@
We first note that $\Delta$ Tot. Bal. $>0\%$ (12m) and $\Delta$ Tot. Bal. $>20\%$ (12m) are binary outcomes which are relatively homoskedastic. $\Delta$ Tot. Bal. (6m) and $\Delta$ Tot. Bal. (12m), measured in Philippine pesos, exhibit more heteroskedasticity. In particular, comparing the treatment group to the marketing group at the 12 month period, we observe a standard deviation ratio of $3.13$. At first glance, this suggests that the FNA might outperform balanced randomization, at least with respect to this specific outcome. This turns out to be false once we investigate other features of $\Delta$ Tot. Bal. (6m) and $\Delta$ Tot. Bal. (12m). Quantiles of these variables are displayed in Table 4. Clearly, they have extremely fat right tails, which we confirm by computing the kurtosis, contained in Table 5. Fat tails worsen the performance of a variety of statistical techniques, including the FNA, as our analysis in Section 4 shows.
Table 3: Heteroskedasticity in Ashraf et al. (2006)
@@
Table 4: Quantiles of Outcome Variables in Ashraf et al. (2006).
@@
Table 5: Kurtosis of the Outcome Variables in Ashraf et al. (2006)
@@
Researchers cannot estimate $C_{m}$ if they only have access to a small pilot. However, this is feasible with data from the full experiment. To do so, we draw samples of size $m$ with replacement and compute $Z_{m}$ for each draw. We then take the average of $Z_{m}+1/Z_{m}$ over 10,000 draws to evaluate $B_{m}$, essentially evaluating the expectation with respect to the empirical CDF. Proposition 4.4 then allows us to compute $C_{m}$. We evaluate $C_{m}$ for $m\leq 2000$, using grids of size $20$. Results are displayed in Figure 5. Given the high kurtosis, there is a relatively large range of ratios of standard deviations for which the balanced allocation is preferred to the FNA.
We can further compute the $m$ necessary before the FNA outperforms the balanced allocation. Given the upper and lower bounds in Figure 5, we smooth over the grid points using cubic interpolation. We then find the minimum $m$ at which the either bound attains the values in Table 3. These are the “Exact" intervals presented in Table 6. Here, we see that the necessary pilot sizes are between 25–50% of the full experiment. For the comparison of the treatment and marketing groups at 6 months, the necessary pilot size exceeds $2,000$, falling outside the set of grid points we explored. To complete the analysis, we use Proposition 4.5 to obtain approximations of the necessary $m$. Specifically, we estimate $V$ using their sample analogs computed on the full experiment:
$\hat{V}=\frac{\frac{1}{n_{1}}\sum_{A_{i}=1}\left(Y_{i}-\overline{Y}(1)\right)^{4}}{\left(\frac{1}{n_{1}-1}\sum_{A_{i}=1}\left(Y_{i}-\overline{Y}(1)\right)^{4}\right)^{2}}+\frac{\frac{1}{n_{0}}\sum_{A_{i}=0}\left(Y_{i}-\overline{Y}(0)\right)^{4}}{\left(\frac{1}{n_{0}-1}\sum_{A_{i}=0}\left(Y_{i}-\overline{Y}(0)\right)^{4}\right)^{2}}-2~{},$
where $n_{a}$ is the number of units with treatment status $a$ in the full experiment and $\overline{Y}(a)=\frac{1}{n_{a}}\sum_{A_{i}=a}Y_{i}$. We can then estimate $m$ by
$\hat{m}=\frac{\hat{V}}{\left(1-\frac{\hat{\sigma}(1)}{\hat{\sigma}(0)}\right)^{2}}~{}.$
Comparing the asymptotic intervals to the exact ones, we see that the former is far too optimistic for the fat-tailed DGP in Ashraf et al. (2006). This is likely because the sub-Gaussian assumption is inappropriate for this data. Nonetheless, applying the asymptotic bounds to the case of the treatment vs marketing groups at 6 months, we find that a pilot size of 7,000 would be necessary for the FNA to outperform balanced randomization. This is nearly 4 times the size of the actual experiment.
@@
In sum, our analysis suggests that the FNA would have performed poorly in the context of Ashraf et al. (2006). Even though the outcomes of interest exhibit stronger heteroskedasticity, the fat tails of the outcome distributions also impedes the estimation of $\widetilde{p}$, so that ultimately, very large pilot sizes are needed for the FNA to outperform the balanced allocation in this example.
@@ 6 Potential Solutions
In this section, we use simulations to explore methods that can improve the performance of the FNA with small pilots. For definitions of the models, see Section 4.4. We consider in turn (1) testing for homoskedasticity, (2) additive regularization and (3) exponential regularization. We conclude with a comparison of these results and speculate that exponential regularization is preferable to the other two methods. Additional simulation results can be found in Appendix C.2.
@@ 6.1 Testing for Homoskedasticity
Since balanced randomization is preferred when outcomes are relatively homoskedastic, consider the following two-step procedure. First test for homoskedasticity in the pilot. If the test rejects, use the FNA in the main wave. Otherwise, use the balanced allocation instead.
@@
$\tilde{W}=\sqrt{\frac{m}{2}}\frac{(\tilde{\sigma}^{2}(1)-\tilde{\sigma}^{2}(0))}{\sqrt{\tilde{V}_{W}(1)+\tilde{V}_{W}(0)}}~{},$
@@
$\tilde{V}_{W}(a)=\begin{pmatrix}1&-2\overline{\widetilde{Y}}(a)\end{pmatrix}\widetilde{\Sigma}_{W}(a)\begin{pmatrix}1\\ -2\overline{\widetilde{Y}}(a)\end{pmatrix}~{},$
$\widetilde{\Sigma}_{W}(a)=\begin{pmatrix}\frac{1}{m/2}\sum_{\widetilde{A}_{i}=a}\left(\widetilde{Y}_{i}^{2}-\overline{\widetilde{Y}^{2}}(a)\right)^{2}&\frac{1}{m/2}\sum_{\widetilde{A}_{i}=a}\left(\widetilde{Y}_{i}-\overline{\widetilde{Y}}(a)\right)\left(\widetilde{Y}_{i}^{2}-\overline{\widetilde{Y}^{2}}(a)\right)\\ *&\frac{1}{m/2}\sum_{A_{i}=a}(\widetilde{Y}_{i}-\overline{\widetilde{Y}}(a))^{2}\end{pmatrix}~{},$
$\overline{\widetilde{Y}}(a)=\frac{1}{m/2}\sum_{A_{i}=a}\widetilde{Y}_{i}\quad,\quad\overline{\widetilde{{Y}}^{2}}(a)=\frac{1}{m/2}\sum_{A_{i}=a}\widetilde{Y}_{i}^{2}~{}.$
@@
$\phi=\mathbf{1}\left\{\tilde{W}>\Phi^{-1}\left(1-\frac{\alpha}{2}\right)\right\}+\mathbf{1}\left\{\tilde{W}<\Phi^{-1}\left(\frac{\alpha}{2}\right)\right\}~{},$
where $\Phi(\cdot)$ is the CDF of the standard normal distribution. Under small pilot asymptotics, this test will not be consistent. However, the two-step procedure will be asymptotically equivalent to the FNA when pilots are large, since for any fixed level $\alpha$, the test will reject with probability approaching $1$ as $m\to\infty$. However, under such a “pre-testing procedure", the main-wave difference-in-means estimator will not be uniformly consistent.
Nonetheless, the procedure appears to perform well in practice. Figure 7 presents the MSEs that obtain using first stage tests with $\alpha\in\left\{1\%,5\%,10\%,20\%,50\%\right\}$. When $\sigma(1)/\sigma(0)\geq 0.5$, decreasing $\alpha$ decreases MSE across all distributions. Conversely, when $\sigma(1)/\sigma(0)\leq 0.5$, decreasing $\alpha$ increases MSE across all distributions. The appropriate level of $\alpha$ then depends on researcher’s preference for the performance of the method over the range of $\sigma(1)/\sigma(0)$.
Figure 6: MSE under testing for homoskedasticity. We compare the Infeasible Neyman Allocation (INA), Balanced Allocation (BA), Feasible Neyman Allocation (FNA), as well as testing for homoskedasticity with various values of $\alpha$ (Test - $\alpha\%$).
Figure 7: MSE under additive regularization. We compare the Infeasible Neyman Allocation (INA), Balanced Allocation (BA), Feasible Neyman Allocation (FNA), as well as additive regularization with various values of $\nu$ (Add - $\nu$).